I'm having a little trouble with this question. I can't find anything in our book that looks anything like it. The instructions are: "solve each equation, inequality or system of equations"
square root x-1=x-1
I tried this
x-1=(x-1)^2 Remove the square root form the left side and square the right.
-x^2+x-1+1^2=0 Set the equation to zero and cancell -1 and +1^2
x^2-x=0
There doesn't seem to be a solution.
Am I even approaching this correctly?

Answer by ankor@dixie-net.com(22740)

square root x-1=x-1
:
x - 1 = (x-1)^2; as you did
:
x - 1 = x^2 - 2x + 1; FOILed
:
0 = x^2 -2x - x + 1 + 1
:
A quadratic equation:
x^2 - 3x + 2 = 0
Factors to
(x-2)(x-1) = 0
:
x = 2, x = 1; either solution will work